OpenStudy (anonymous):

x - 0.2x + 0.06(x - 0.2x) = 66.91 help!

OpenStudy (anonymous):

Someone, anyone please respond !

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

x - 0.2x + 0.06(x - 0.2x) = 66.91 0.8x + 0.06x - 0.012x = 66.91 0.848x=66.91 x=66.91/0.848

OpenStudy (anonymous):

x = 78.9

OpenStudy (anonymous):

that is not one of the answer choices

OpenStudy (anonymous):

thx